OpenStudy (idku):

Checking work on changing the boundaries of a definite integral (I have been taught to sub back, but going over this approach now)

OpenStudy (idku):

\[2\int\limits_{4}^{11}x(x^2-15)^3~dx\]

OpenStudy (idku):

\[\color{blue}{u=x^2-15}\]\[x=4,~~~~~~~\Rightarrow~~~~~~~u=4^2-15~~~~~~~\Rightarrow~~~~~~~u=1\]\[x=11,~~~~~\Rightarrow~~~~~u=11^2-15~~~~~\Rightarrow~~~~~u=121-15~~~~~\Rightarrow~~~~~u=106\]\[\color{blue}{du=2x~dx}\] \[\int\limits_{1}^{106}u^3~du\]

OpenStudy (idku):

So, now all what I have to do is to integrate it (without subbing the x back into the equation. \[=\frac{1}{4}(106)^4-(1)^4\]and then just a calculation.

OpenStudy (idku):

@whpalmer4 @triciaal can you guys check my work please?

OpenStudy (idku):

oh 1/4 next to (1)^4.

OpenStudy (idku):

but anything other than this?

OpenStudy (whpalmer4):

appears to get the right answer

OpenStudy (idku):

(Wouldn't use geometric shapes on this one, whether on the one with u's or x's) ty